A thin non-conducting rod of finite length L has a charge q spread uniformly along it. Show that the magnitude of the electric field at point P a distance y on the perpendicular bisector of the rod is given by: E = q 2π0y 1 (L2 + 4y 2) 1/2


1
Expert's answer
2021-07-05T08:56:34-0400


q= net charge

Liner charge density

"λ= \frac{q}{L}"

L=length of rod

dq=λdx

Electric field due to dq as p

"dE = \frac{kdq}{γ^2} \\

γ^2 = x^2 + y^2 \\

E_x = \int dEsinθ = \int \frac{kdq}{γ^2} \times \frac{x}{γ} \\

= \int \frac{k \times λ \times dx \times x}{(x^2+y^2)^{3/2}} \\

= k λ \int^{L/2}_{-L/2} \frac{xdx}{(x^2+y^2)^{3/2}} \\

E_x=0 \\

E_y= \int dEcosθ = \int \frac{kdq}{γ^2} \times \frac{y}{γ} \\

= \int \frac{k \times λ \times dx \times y}{γ^3} \\

= kλ y \int^{L/2}_{-L/2} \frac{dx}{(x^2+y^2)^{3/2}}"